If a man is n years old in the year n-square then what is his age in 1960?
Answer by Edwin McCravy(20055)

The way you have it worded with the present tense verb "is" makes it confusing
because this year is 2014, many years after 1960. It should be "What WAS
(not "is") his age in 1960?" 

Let's find the two square years that 1960 is between.

√1960 = 44.27188724, which is between 44 and 45.

Therefore 1960 is between the two square years 44² = 1936 and 45² = 2025.

So he was either born in the year 1936-44 = 1892 or the year 2025-45 = 1980.

He couldn't have been born in 1980 since 1960 was 20 years before he was born!

So he was born in 1892 (and is now dead or else 122 years old).  

Since 1960 was 1960-1892=68 years after 1892, he was 68 in 1960.

[  Incidentally what you can learn from this problem is that if you know
somebody who was born in 1980, you can tell them that they will be n in the
year n².  Such a person would now be 2014-1980 = 34 years old.  Do you know
anybody who is 34 now in 2014?  If so, you can have fun telling them that. :)  ]

Edwin